OpenStudy (anonymous):

Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation. sine of x divided by one minus cosine of x + sine of x divided by one minus cosine of x = 2 csc x

OpenStudy (anonymous):

\[(\sin x/1 - \cos x) + (\sin x/1 + \cos x) = 2 \csc x\]

OpenStudy (anonymous):

@AnswerMyQuestions help?

OpenStudy (anonymous):

@CrazyMexican816 help?

OpenStudy (anonymous):

@ilovebmth1234

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@NerdyBookWorm

OpenStudy (anonymous):

ok,

OpenStudy (anonymous):

sorry im trying to figure this out on paper lol one sec

OpenStudy (anonymous):

ok, d\[sinx/1-cosx + sinx/1+cosx = 2cosx\]oes your equation looks similar to this?

OpenStudy (anonymous):

2cscx *

OpenStudy (anonymous):

Yeah :)

OpenStudy (anonymous):

ok well @iGreen

OpenStudy (anonymous):

im kinda confused

OpenStudy (anonymous):

Any help?

OpenStudy (anonymous):

You there?

OpenStudy (anonymous):

Can you make the denominators equal? Or Do you know how to do that?

OpenStudy (anonymous):

I have no idea how to do that

OpenStudy (anonymous):

Okay, It is a method called LCM : Least Common Multiple..

OpenStudy (anonymous):

But, I will not go deep into that, because you may confuse, I have one more simpler idea..

OpenStudy (anonymous):

\[\frac{\sin(x)}{1 - \cos(x)} \times \frac{1 + \cos(x)}{1 + \cos(x)} + \frac{\sin(x)}{1 + \cos(x)} \times \frac{1 - \cos(x)}{1 - \cos(x)}\] Can I do like this?

OpenStudy (anonymous):

Yeah I assume so

OpenStudy (anonymous):

I have multiplied and divided by the same quantity, so it will be cancelled, giving out the same result as we had earlier.. :)

OpenStudy (anonymous):

So, I also assume that.. :P

OpenStudy (anonymous):

You are not supposed to assume, just look at it and find out why I have written that only.. In other words, in your mind it should strike that this is also a "Rationalization Process".. :)

OpenStudy (anonymous):

There, as denominators are same, you can just add the numerator terms.. :)

OpenStudy (anonymous):

\[\frac{\sin(x) - \sin(x)\cos(x) + \sin(x) + \sin(x)\cos(x)}{(1-\cos(x)(1+\cos(x))} \implies \frac{2\sin(x)}{1 - \cos^2(x)} \implies \frac{2\sin(x)}{\sin^2(x)}\]

OpenStudy (anonymous):

Next step I leave it on you.. :P

OpenStudy (anonymous):

I'm so lost. I get what you're saying about the denominators being the same. But I really don't get this.

OpenStudy (anonymous):

May I ask something? You have two persons in your pic, which one is you??

OpenStudy (anonymous):

The girl on the left

OpenStudy (anonymous):

Type the thing where you have lost.. :)

OpenStudy (anonymous):

I was thinking the same, as you have said, I am lost.. :P

OpenStudy (anonymous):

See: Look here: Suppose you have fraction like 5/3.. Now if I multiply it and divide it by same number let us say 4, then our fraction remains same: \(\frac{5}{3} \times \frac{4}{4}\) = \(\frac{5}{3}\) This you are getting?

OpenStudy (anonymous):

Yeah I get that

OpenStudy (anonymous):

Look above carefully, I have done the same there..

OpenStudy (anonymous):

I have multiplied and divide by the same quantity.. Yes??

OpenStudy (mathmath333):

\(\large\tt \color{black}{\dfrac{sinx}{1-cosx}+\dfrac{sinx}{1+cosx}}\) \(\large\tt \color{black}{=sinx(\dfrac{1}{1-cosx}+\dfrac{1}{1+cosx})}\) \(\large\tt \color{black}{=sinx(\dfrac{1+cosx+1-cosx}{1-cos^2x})}\) \(\large\tt \color{black}{=sinx(\dfrac{2}{sin^2x})}\) \(\large\tt \color{black}{=(\dfrac{2}{sinx})}\) \(\large\tt \color{black}{=2cscx}\)

OpenStudy (anonymous):

@mathmath333 again it comes to making denominators equal, and that is the part she is not getting.. :P

OpenStudy (anonymous):

@mathmath333 you're a lifesaver. That makes so much sense. Okay thanks :)

OpenStudy (anonymous):

Okay my mistake, it was not the denominator part, it was the whole solution which was making her lost.. :P

OpenStudy (mathmath333):

well i did the same thing as waterinyes i just took the denominator out